OpenStudy (anonymous):

PLZ HELP...I GIVE OUT MEDALS!!!!! Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). A.) y = 4/3x + 8 B.) y = 4/3x + 3 C.) y = −3/4x + 8 D.) y = −3/4x + 3

hartnn (hartnn):

can you find slope of the line 4x − 3y = −12

OpenStudy (haseeb96):

hartnn I can find it

hartnn (hartnn):

i know :) but its turtle's question let her try ...

OpenStudy (haseeb96):

give me some clue to solve this questio n

OpenStudy (anonymous):

i don't know how to find the slope...

hartnn (hartnn):

try to isolate 'y' from 4x − 3y = −12 to bring it in the form of y=mx+c

OpenStudy (anonymous):

so... if I get rid of y in the equation, it would be 4x - 3 =-12 right...? then if I work the problem like that I could probably figure out what X is...am I correct?

OpenStudy (haseeb96):

-3y=-12-4x y=12+4x\3 y=4 + 4x\3

hartnn (hartnn):

you can't just drop the 'y'!

OpenStudy (haseeb96):

hartnn help me

OpenStudy (anonymous):

I know that...then when I find X...I can put Y back in then figure out what that is then go back and finish the problem...right?

hartnn (hartnn):

no... we are just rearranging the terms here... 4x-3y = -12 subtract 4x from both sides, what u get ?

OpenStudy (anonymous):

-3y = -8

hartnn (hartnn):

no you cannot combine -12 and 4x or just drop the x you would get -3y = -4x -12 got this ?

OpenStudy (anonymous):

oh...I forgot you could only do that if they had the same letter behind it...or no letter at all

hartnn (hartnn):

thats correct

OpenStudy (anonymous):

soo....would it be A?

hartnn (hartnn):

no.... you got this ? -3y = -4x -12 the parallel line equation will be of the form -3y = -4x +c you just need to find value of c